OpenStudy (anonymous):

N+3/4=19/20. What is N = to?

OpenStudy (anonymous):

(19/20)-(3/4)= 1/5, n= 1/5

OpenStudy (anonymous):

Thanks! so then what would N- 1/5= 7/20 be?

OpenStudy (anonymous):

(7/20)+(1/5)= 11/20

OpenStudy (anonymous):

Wait I still dont get how that ='s 11/20 cuz, 7+1=8... and 20 and 20= 20

OpenStudy (anonymous):

when u add fractions, u cant really add across until u find a common denominator. so (7/20)+(1/5), to get a common denominator, u can just multiply 5*4 to get 20, then u would have to multiply 4 with the number at the top, which is 1. so u get (7/20)+(4/20), u have a common denominator now, so u can just add across, 7+4=11, and u dont ever really add denominators together bc they are the same

OpenStudy (anonymous):

Thanks so much!!

OpenStudy (anonymous):

np :) hope my explanation helps

OpenStudy (anonymous):

:), Yes it does! You do it the old fashion way, when a negative number turns into a positive and, negative turns into postive. You see i have a math test tomorrow so I wanna see if i was doing it right! hahaa

OpenStudy (anonymous):

wo